1.A factory produced 593 257 vaccines a week. However, 5 728 of them were damaged. How many were good?
2.The price of gasoline originally cost ₱ 65.00. If the price was marked up by 25%, how much was the new price?
3.Mr. Santos divided 3 542 square meters of a lot into 23 equal parts. How large was each lot?
4.A piece of wood is 24 3/8 inches long. How many 4 7/8 inches pieces of wood can be cut from it?
5.Joan’s mother earned ₱ 556, ₱ 654, and ₱ 482 on Monday, Tuesday, and Wednesday respectively in her sari - sari store. How much is her average sale in three days?
pls with solution

Answer by Boreal(15235)

1. It is 593257-5728=587,529
2. The markup is 0.25*65.00 or (1/4)*65=$16.25 (16 1/4) so the new price is $81.25
3. this is 3542/23=154 by dividing. And 154*23=154*20+154*3=3080+462=3542 to check
4.24 3/8 is almost 25 and 4 7/8 is almost 5, so I would start with 5 pieces
5*4 7/8=20+35/8, and 35/8 is 4 3/8 so 5 equal pieces are indeed 24 3/8 inches
5. This is the sum divided by 3 or 1692/3=564 P.
Check the differences of each from 564 and their sum should be 0.
-8+90-82, and their sum is 0.